1.Factor using difference of cubes (a-b)(a2+ab+b2)

8x3-27


2.Factor by grouping 2x3+7x2+2x+7


3.List the possible roots using the Rational Root Theorem x= +- p/q where q is the leading coefficient and p is the constant

f(x) = 3x5 - 2x2 - 7x + 4

1
Expert's answer
2021-05-19T16:08:35-0400

"1) 8x^3-27=(2x-3)(4x^2)+6x+9)"

"2) 2x^3+7x^2-7x+4=x(x^2+1)+7(x^2+1)=(2x+7)(x^2+1)"

"3) 3x^5-2x^2-7x+4=3m^5\/n^5-2m^2\/n^2-7m\/n+4=0"

"3m^5=2m^2n^3+7mn^4-4n^5"

"3m^5=n^3(2m^2+7mn-4n^2)"

"3m^5\/n^3" divides  entirely

n can be 1, 3, -1 ,-3

(possible denominators of the result)

"3m^5-2m^2n^3-7mn^4=-4n^5"

"m(-3m^4+2mn^3+7n^4)=4n^5"

"4n^5\/m" divides entirely

m can be 1, 4, -1, -4, 2, -2

(possible numerators of the result)

m/n (or p/q) can be the one of combinations of this numbers:

1, -1, 1/3, -1/3, 4, -4, 4/3, -4/3, 2, -2, 2/3, -2/3

Using the substitution method with each of these possiblle roots,

we can find out that m/n=x=2/3.

It is the root of this polynom at which it vanishes, takes a zero value.